OpenStudy (anonymous):

A line has a rise of 6 and a slope of 1/20. What is the run? 30 90 120

OpenStudy (anonymous):

You just want the answer or the method?? Because I'll help you with the method! And you will hav to find it!!

OpenStudy (anonymous):

just tell me how to solve it times divide whatever

OpenStudy (anonymous):

Slope=Rise/Run Use the formula Substitute the Values Get ur answer.....Cheers!!:)

OpenStudy (anonymous):

1/20=6/30

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so a would be my awnser

OpenStudy (anonymous):

@amistre64 ....@ivettef365

OpenStudy (anonymous):

U formula substitution is wrong!!

OpenStudy (anonymous):

it shud be

OpenStudy (amistre64):

\[\frac{rise}{run}:\frac 6x=\frac 1{20}\] x = 6(20)

OpenStudy (anonymous):

1/20=6/x solve for x

OpenStudy (anonymous):

i already did omg i said 30

OpenStudy (anonymous):

How do u get 30 if u solve for x??

OpenStudy (anonymous):

On solving the above equality we get x=6*20

OpenStudy (anonymous):

120